# Robust Decision Transformer: Tackling Data Corruption in Offline RL via Sequence Modeling

### v1

@@ 4.4 Ablation Study
We conduct comprehensive ablation studies to analyze the impact of each component on RDT’s robustness. For evaluation, we use the "walker2d-medium-replay", "kitchen-compete", and "relocate-exper"’ datasets. Specifically, we compare several variants of RDT: (1) DT(RP), which incorporates reward prediction in addition to the original DT; (2) DT(RP) w. ED, which adds embedding dropout to DT(RP); (3) DT(RP) w. GWL, which applies Gaussian weighted learning on top of DT(RP); and (4) DT(RP) w. IDC, which integrates only the iterative data correction method.
We evaluate the performance of these variants under different data corruption scenarios, as depicted in Figure 5. In summary, all variants demonstrate improvements over DT(RP), proving the effectiveness of the individual components. Notably, Gaussian weighted learning appears to provide the most significant contribution, particularly under reward attack. However, it is important to note that none of these tailored models outperform RDT, indicating that the integration of all proposed techniques is crucial for achieving optimal robustness. Additionally, we provide ablation results on reward prediction in Section D.2, which demonstrate that reward prediction also brings improvements under state and action corruption compared to DT.
Figure 5: Ablation study on the impact of different proposed components.
In this section, we further investigate the robustness of RDT when deployed in perturbed environments after being trained on corrupted data, a challenging setting that includes both training-time and testing-time attacks. To address this, we evaluate RDT under two types of observation perturbations during the testing phase: random and action diff, following prior works (Yang et al., 2022a, ; Zhang et al.,, 2020). The perturbation scale is used to control the extent of influence on the observation. Detailed implementation of observation perturbations during the testing phase is provided in Section C.2.
The comparison results are presented in Figure 6, where all algorithms are trained on the offline dataset with action corruption and evaluated under observation perturbation. These results demonstrate the superior robustness of RDT under the two types of observation perturbations, maintaining stability even at a high perturbation scale of 0.5. Notably, RORL (Yang et al., 2022a, ), a SOTA offline RL method designed to tackle testing-time observation perturbation, fails to perform effectively on these tasks. Additionally, DT and RIQL experience significant performance drops as the perturbation scale increases. Further results comparing different algorithms trained on offline datasets with state or reward corruption can be found in Section D.4.
(a) Random observation perturbation.
(b) "Action diff" observation perturbation.
Figure 6: Performance under various observation perturbation scales during the testing phase. All the algorithms are trained under random action corruption during the training phase.
@@ 5 Related Work
In offline RL, several works have focused on testing-time robustness against environment shifts (Shi and Chi,, 2022; Yang et al., 2022a, ; Panaganti et al.,, 2022; Zhihe and Xu,, 2023). Regarding training-time robustness, Li et al., (2023) explore various types of reward attacks in offline RL and find that certain biases can inadvertently enhance the robustness of offline RL methods to reward corruption. From a theoretical perspective, Zhang et al., (2022) propose a robust offline RL algorithm utilizing robust supervised learning oracles. Additionally, Ye et al., 2024b employ uncertainty weighting to address reward and dynamics corruption, providing theoretical guarantees. The most relevant research by Yang et al., 2024b employs the Huber loss to handle heavy-tailedness and utilizes quantile estimators to balance penalization for corrupted data. Furthermore, Mandal et al., (2024); Liang et al., (2024) enhance the resilience of offline algorithms within the RLHF framework. It is important to note that these studies primarily focus on enhancing temporal difference methods, with no emphasis on leveraging sequence modeling techniques to tackle data corruption.
@@
Recent research has redefined offline RL decision-making as a sequence modeling problem using Transformer architectures (Chen et al.,, 2021; Janner et al.,, 2021). Unlike traditional RL methods, these studies treat RL as a supervised learning task at a trajectory level. A seminal work, Decision Transformer (DT) (Chen et al.,, 2021), uses trajectory sequences to predict subsequent actions. Trajectory Transformer (Janner et al.,, 2021) discretizes input sequences into tokens and employs beam search to predict the next action. These efforts have led to subsequent advancements. For instance, Prompt DT (Xu et al.,, 2022) integrates demonstrations for better generalization, while Xie et al., (2023) introduce pre-training with future trajectory information. Q-learning DT (Yamagata et al.,, 2023) refines return-to-go using Q-values, and Agentic Transformer (Liu and Abbeel,, 2023) uses hindsight to relabel target returns. LaMo (Shi et al.,, 2023) leverages pre-trained language models for offline RL, and DeFog (Hu et al.,, 2022) addresses robustness in specific frame-dropping scenarios. Our work deviates from these approaches by focusing on improving robustness against data corruption in offline RL.
@@ 6 Conclusion
In this study, we investigate the robustness of offline RL algorithms under various data corruptions, with a specific focus on sequence modeling methods. Our empirical evidence suggests that current offline RL algorithms based on temporal difference learning exhibit significant susceptibility to data corruption, particularly in scenarios with limited data. To address this issue, we introduce the Robust Decision Transformer (RDT), a novel robust offline RL algorithm developed from the perspective of sequence modeling. Our comprehensive experiments highlight RDT’s exceptional robustness against both random and adversarial data corruption, across different corruption ratios and scales. Furthermore, we demonstrate RDT’s superiority in handling both training-time and testing-time attacks. We hope that our findings will inspire further research into the exploration of using sequence modeling methods to address data corruption challenges in increasingly complex and realistic scenarios.
@@ Appendix A Algorithm Pseudocode
To provide an overview and better understanding, we detail the implementation of Robust Decision Transformer (RDT) in  Algorithm 1.
@@
6:     Update corresponding mean $\mu_{\delta}$ and variance $\sigma^{2}_{\delta}$
@@
10:     Substitute corrupted actions and rewards with the predicted actions and rewards in dataset $\mathcal{D}$.
@@ Appendix B Additional Related Works
Offline RL. Maintaining proximity between the policy and data distribution is essential in offline RL, as distributional shifts can lead to erroneous estimations (Levine et al.,, 2020). To counter this, offline RL algorithms are primarily divided into two categories. The first category focuses on policy constraints on the learned policy (Wang et al.,, 2018; Fujimoto et al.,, 2019; Li et al.,, 2020; Fujimoto and Gu,, 2021; Kostrikov et al.,, 2021; Emmons et al.,, 2021; Yang et al., 2022b, ; Sun et al.,, 2024; Xu et al.,, 2023; Park et al.,, 2024). The other category learns pessimistic value functions to penalize OOD actions (Kumar et al.,, 2020; Yu et al.,, 2020; An et al.,, 2021; Bai et al.,, 2022; Yang et al., 2022a, ; Ghasemipour et al.,, 2022; Sun et al.,, 2022; Nikulin et al.,, 2023; Huang et al.,, 2024). To enhance the potential of offline RL in handling more complex tasks, recent research has integrated advanced techniques like GAN (Vuong et al.,, 2022; Wang et al.,, 2023), transformers (Chen et al.,, 2021; Chebotar et al.,, 2023; Yamagata et al.,, 2023) and diffusion models (Janner et al.,, 2022; Hansen-Estruch et al.,, 2023; Wang et al.,, 2022).
@@ C.1 Data Corruption Details during Training Phase
Our study utilizes both random and adversarial corruption across three elements: states, actions, and rewards. We consider a range of tasks including MuJoCo, Kitchen, and Adroit. Particularly, we utilize the “medium-replay-v2” datasets in the MuJoCo tasks with sampling ratios of 100% and 10%, the “expert-v0” datasets in the Adroit tasks with a sampling ratio of 1%, and we employ full datasets for the tasks in the Kitchen due to their already limited data size. These datasets (Fu et al.,, 2020) are collected either during the training process of an SAC agent or from expert demonstrations, thereby providing highly diverse and representative tasks of the real world. To control the overall level of corruption within the datasets, we introduce two parameters $c$ and $\epsilon$ following previous work (Yang et al., 2024b, ). The parameter $c$ signifies the rate of corrupted data within a dataset, whilst $\epsilon$ represents the scale of corruption observed across each dimension. We outline three types of random data corruption and present a comprehensive overview of a mixed corruption approach as follows. Note that in our setting, only three independent elements (i.e., states, actions, and rewards) are considered under the trajectory-based storage approach.
@@
  • Random reward attack: We randomly sample $c\cdot N\cdot T$ rewards from all given trajectories, and modify the reward to $\hat{r}\sim\text{Uniform}[-30\cdot\epsilon,30\cdot\epsilon]$. We multiply by 30 because we have noticed that offline RL algorithms tend to be resilient to small-scale random reward corruption (also observed in (Li et al.,, 2023)), but would fail when faced with large-scale random reward corruption.
@@
  • Adversarial state attack: We first pretrain IQL agents with a Q function $Q_{p}$ and policy function $\pi_{p}$ on clean datasets. Then, we randomly sample $c\cdot N\cdot T$ states, and modify them to $\hat{s}=\min_{\hat{s}\in\mathbb{B}_{d}(s,\epsilon)}Q_{p}(\hat{s},a)$. Here, $\mathbb{B}_{d}(s,\epsilon)=\{\hat{s}||\hat{s}-s|\leq\epsilon\cdot\text{std}(s)\}$ regularizes the maximum difference for each state dimension. The optimization is implemented through Projected Gradient Descent similar to prior works (Madry et al.,, 2017; Zhang et al.,, 2020; Yang et al., 2024b, ). Specifically, We first initialize a learnable vector $z\in[-\epsilon,\epsilon]^{d_{s}}$, and then conduct a 100-step gradient descent with a step size of 0.01 for $\hat{s}=s+z\cdot\text{std}(s)$, and clip each dimension of $z$ within the range $[-\epsilon,\epsilon]$ after each update.
@@ C.2 Observation Perturbation Details during the Testing Phase
We evaluate the Robust Decision Transformer (RDT) under two types of perturbations during the testing phase: random and action diff perturbations, as described in prior works (Yang et al., 2022a, ; Zhang et al.,, 2020). Both perturbation types target the observations during the testing phase. The detailed implementations are as follows:
  • Random: We sample perturbed states within an $l_{\infty}$ ball of norm $\epsilon$. Specifically, we create the perturbation set $\mathbb{B}_{d}(s,\epsilon)=\{\hat{s}:d(s,\hat{s})\leq\epsilon\}$ for state $s$, where $d(\cdot)$ is the $l_{\infty}$ norm, and sample one perturbed state to return to the agent.
  • Action Diff: This is an adversarial attack based on the pretrained IQL deterministic policy $\mu(s)$. We first sample 50 perturbed states within an $l_{\infty}$ ball of norm $\epsilon$ and then find the one that maximizes the difference in actions: $\max_{\hat{s}\in\mathbb{B}_{d}(s,\epsilon)}||\mu(s)-\mu(\hat{s})||^{2}$.
The parameter $\epsilon$ controls the extent of the observation perturbation. In this experiment, we first train RDT under different data corruption scenarios and then evaluate its performance in environments subjected to observation perturbation.
@@ C.3 Implementation Details of RDT
We implement the DT and RDT algorithms using the existing code base***https://github.com/tinkoff-ai/CORL. Specifically. we build the network with 3 Transformer blocks, incorporating one MLP embedding layer for each key element: state, action, and return-to-go. We update the neural network using the AdamW optimizer, with a learning rate set at $1\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. The batch size is set to 64, with a sequence length of 20, to ensure effective and efficient training. To maintain stability during training, we adopt the state normalization as in (Yang et al., 2024b, ). During the training phase, we train DeFog, DT and RDT for 100 epochs and other baselines with 1000 epochs. Each epoch is characterized by 1000 gradient steps. For evaluative purposes, we rollout each agent in the clean environment across 10 trajectories, each with a maximum length of 1000, and we average the returns from these rollouts for comparison. We ensure the consistency and reliability of our reported results by averaging them over 4 unique random seeds. All experiments are conducted on 8 P40 GPUs.
As for the hyperparameters of RDT, we configure the values for the embedding dropout probability as $p$, Gaussian Weighted coefficient $\beta_{a}$ and $\beta_{r}$, and iterative data correction threshold $\zeta$ through the sweeping mechanism. We start the iterative data correction at the 50th epoch for all tasks. We use iterative action correction exclusively for action corruption and iterative reward correction exclusively for reward corruption; therefore, only a single $\zeta$ value is shown in Table 3. It’s noteworthy that we do not conduct a hyperparameter search for individual datasets; instead, we select hyperparameters that apply to the entire task group. As such, RDT doesn’t require extensive hyperparameter tuning as previous methods. The exact hyperparameters used for the random corruption experiment are detailed in Table 3. We employ the same hyperparameters for the adversarial corruption experiment, further attesting to the robustness and stability of RDT.
Table 3: Hyperparameters used for RDT under the random corruption benchmark.
@@
MuJoCo state 0.0 (1.0, 10.0) --
action 0.0 (0.1, 0.0) 3.0
reward 0.0 (0.1, 0.1) 6.0
MuJoCo (10%) state 0.1 (0.1, 0.1) --
action 0.1 (0.1, 0.1) 5.0
reward 0.2 (0.1, 0.1) 6.0
Kitchen state 0.1 (10.0, 10.0) --
action 0.1 (10.0, 10.0) 1.5
reward 0.0 (10.0, 10.0) 3.0

### v3

@@ 4.4 Ablation Study
We conduct comprehensive ablation studies to analyze the impact of each component on RDT’s robustness. For evaluation, we use the “walker2d-medium-replay”, “kitchen-compete”, and “relocate-expert” datasets. Specifically, we compare several variants of RDT: (1) DT(RP), which incorporates reward prediction in addition to the original DT; (2) DT(RP) w. ED, which adds embedding dropout to DT(RP); (3) DT(RP) w. GWL, which applies Gaussian weighted learning on top of DT(RP); and (4) DT(RP) w. IDC, which integrates only the iterative data correction method.
We evaluate the performance of these variants under different data corruption scenarios, as depicted in Figure 6. In summary, all variants demonstrate improvements over DT(RP), proving the effectiveness of the individual components. Notably, Gaussian weighted learning appears to provide the most significant contribution, particularly under reward attack. However, it is important to note that none of these tailored models outperform RDT, indicating that the integration of all proposed techniques is crucial for achieving optimal robustness. Additionally, while reward prediction could potentially lead to performance drops under reward attacks, it brings obvious improvements under action and state attacks for specific tasks. More ablation results about three individual component and reward prediction are provided in Section D.9 and Section D.2.
Figure 6: Ablation study on the impact of different proposed components of RDT.
4.5 Impact of Transformer Backbone on Baselines
To study the impact of the transformer backbone on baselines, we adapted temporal difference learning baselines such as UWMSG and RIQL by incorporating the same transformer backbone used in DT and RDT. Specifically, we modified UWMSG and RIQL by equipping their policy networks with the transformer backbone while keeping the value networks unchanged. These modified algorithms are referred to as UWMSG w. TB and RIQL w. TB.
Figure 7: Ablation study on the impact of transformer backbone on baselines.
As shown in Figure 7, while the transformer backbone does not improve UWMSG, it significantly enhances RIQL’s overall performance, particularly in the challenging Adroit task. The ineffectiveness of the sequence model on UWMSG is likely due to the inherent difficulties of Q-value learning. These comparative results suggest that merely adopting a transformer structure does not directly enable these baselines to outperform RDT; in some cases, it may even decrease performance.
@@ 5 Related Work
Several works have focused on testing-time robustness against environment shifts (Shi & Chi, 2022; Yang et al., 2022a; Panaganti et al., 2022; YANG & Xu, 2024). Regarding training-time robustness, Li et al. (2023) explores various types of reward attacks in offline RL and finds that certain biases can inadvertently enhance the robustness of offline RL methods to reward corruption. From a theoretical perspective, Zhang et al. (2022) propose a robust offline RL algorithm utilizing robust supervised learning oracles. Ye et al. (2024b) employ uncertainty weighting to address reward and dynamics corruption, providing theoretical guarantees. Ackermann et al. (2024) investigates a similar setting, focusing on trajectory-level non-stationarity in both rewards and dynamics. The most relevant research by Yang et al. (2024b) employs the Huber loss to handle heavy-tailedness and utilizes quantile estimators to balance penalization for corrupted data. Mandal et al. (2024); Liang et al. (2024) enhance the resilience of offline algorithms within the RLHF framework. It is important to note that these studies primarily focus on enhancing temporal difference methods, with no emphasis on leveraging sequence modeling techniques to tackle data corruption.
@@
Recent research has redefined offline RL decision-making as a sequence modeling problem using Transformer architectures (Chen et al., 2021; Janner et al., 2021). Unlike traditional RL methods, these studies treat RL as a supervised learning task at a trajectory level. A seminal work, Decision Transformer (DT) (Chen et al., 2021), uses trajectory sequences to predict subsequent actions. Trajectory Transformer (Janner et al., 2021) discretizes input sequences into tokens and employs beam search to predict the next action. These efforts have led to subsequent advancements. For instance, Prompt DT (Xu et al., 2022) integrates demonstrations for better generalization, while Xie et al. (2023) introduces pre-training with future trajectory information. Q-learning DT (Yamagata et al., 2023) refines the return-to-go using Q-values, while the Agentic Transformer (Liu & Abbeel, 2023) employs hindsight to relabel target returns. Additionally, several other works (Hu et al., 2024; Zhuang et al., 2024) utilize learned value functions to improve policy optimization. LaMo (Shi et al., 2023) leverages pre-trained language models for offline RL, and DeFog (Hu et al., 2022) addresses robustness in specific frame-dropping scenarios. Our work deviates from these approaches by focusing on improving robustness against data corruption in offline RL.
@@ 6 Conclusion
In this study, we investigate the robustness of offline RL algorithms under various data corruption scenarios, with a specific focus on sequence modeling methods. Our empirical evidence suggests that current offline RL algorithms based on temporal difference learning are significantly susceptible to data corruption, especially in scenarios with limited data. To address this issue, we introduce the Robust Decision Transformer (RDT), a novel robust offline RL algorithm developed from the perspective of sequence modeling. Our comprehensive experiments highlight RDT’s excellent robustness against various types of data corruption. Furthermore, we demonstrate RDT’s superiority in handling both training-time and testing-time attacks. We hope that our findings will inspire further research into using sequence modeling methods to address data corruption challenges in increasingly complex and realistic scenarios.
Ethics Statement:
This research was conducted in compliance with all applicable ethical guidelines and institutional regulations. Since the study did not involve human participants, animals, or sensitive data, no specific ethical approvals were required. All data used in this research were obtained from publicly available sources, ensuring full transparency and reproducibility of the results.
Reproducibility Statement:
We have taken several measures to ensure the reproducibility of our results. All datasets used in our experiments are publicly available and can be accessed through https://github.com/Farama-Foundation/D4RL and https://github.com/polixir/NeoRL. The implementation of all baselines and RDT is based on the codebase available at https://github.com/tinkoff-ai/CORL. Detailed instructions for setting up the environment and running the experiments are provided in Appendix C. Our implementation code is available at https://github.com/jiawei415/RobustDecisionTransformer. We encourage researchers to refer to Appendix C for more detailed information.
@@ Appendix A Algorithm Pseudocode
To provide an overview and better understanding, we detail the implementation of the Robust Decision Transformer (RDT) in  Algorithm 1.
@@
6:     Update corresponding mean $\mu_{\delta}$ and variance $\sigma^{2}_{\delta}$.
@@
10:     Substitute actions and rewards with the predicted actions and rewards in dataset $\mathcal{D}$.
@@ Appendix B Additional Related Works
Offline RL. Maintaining proximity between the policy and data distribution is essential in offline RL, as distributional shifts can lead to erroneous estimations (Levine et al., 2020). To counter this, offline RL algorithms are primarily divided into two categories. The first category focuses on policy constraints on the learned policy (Wang et al., 2018; Fujimoto et al., 2019; Li et al., 2020; Fujimoto & Gu, 2021; Kostrikov et al., 2021; Emmons et al., 2021; Yang et al., 2022b; 2023; Sun et al., 2024; Xu et al., 2023; Park et al., 2024). The other category learns pessimistic value functions to penalize OOD actions (Kumar et al., 2020; Yu et al., 2020; An et al., 2021; Bai et al., 2022; Yang et al., 2022a; Ghasemipour et al., 2022; Sun et al., 2022; Nikulin et al., 2023; Huang et al., 2024). To enhance the potential of offline RL in handling more complex tasks, recent research has integrated advanced techniques like GAN (Vuong et al., 2022; Wang et al., 2023), transformers (Chen et al., 2021; Chebotar et al., 2023; Yamagata et al., 2023) and diffusion models (Janner et al., 2022; Hansen-Estruch et al., 2023; Wang et al., 2022).
@@ C.1 Data Corruption Details during Training Phase
Our study utilizes both random and adversarial corruption across three elements: states, actions, and rewards. We consider a range of tasks including MuJoCo, Kitchen, and Adroit. Particularly, we utilize the “medium-replay-v2” datasets in the MuJoCo tasks with sampling ratios of 10%, the “expert-v0” datasets in the Adroit tasks with a sampling ratio of 1%, and we employ full datasets for the tasks in the Kitchen due to their already limited data size. These datasets (Fu et al., 2020) are collected either during the training process of an SAC agent or from expert demonstrations, thereby providing highly diverse and representative tasks of the real world. To control the overall level of corruption within the datasets, we introduce two parameters $c$ and $\epsilon$ following previous work (Ye et al., 2024b; Yang et al., 2024b). The parameter $c$ signifies the rate of corrupted data within a dataset, whilst $\epsilon$ represents the scale of corruption observed across each dimension. We outline three types of random data corruption and present a comprehensive overview of a mixed corruption approach as follows. Note that in our setting, only three independent elements (i.e., states, actions, and rewards) are considered under the trajectory-based storage approach.
@@
  • Random reward attack: We randomly sample $c\cdot N\cdot T$ rewards from all given trajectories, and modify the reward to $\hat{r}\sim\text{Uniform}[-30\cdot\epsilon,30\cdot\epsilon]$. We multiply by 30 because we have noticed that offline RL algorithms tend to be resilient to small-scale random reward corruption (also observed in Li et al. (2023)), but would fail when faced with large-scale random reward corruption.
@@
  • Adversarial state attack: We first pretrain IQL agents with a Q function $Q_{p}$ and policy function $\pi_{p}$ on clean datasets. Then, we randomly sample $c\cdot N\cdot T$ states, and modify them to $\hat{s}=\min_{\hat{s}\in\mathbb{B}_{d}(s,\epsilon)}Q_{p}(\hat{s},a)$. Here, $\mathbb{B}_{d}(s,\epsilon)=\{\hat{s}||\hat{s}-s|\leq\epsilon\cdot\text{std}(s)\}$ regularizes the maximum difference for each state dimension. The optimization is implemented through Projected Gradient Descent similar to prior works (Madry et al., 2017; Zhang et al., 2020; Yang et al., 2024b). Specifically, We first initialize a learnable vector $z\in[-\epsilon,\epsilon]^{d_{s}}$, and then conduct a 100-step gradient descent with a step size of 0.01 for $\hat{s}=s+z\cdot\text{std}(s)$, and clip each dimension of $z$ within the range $[-\epsilon,\epsilon]$ after each update.
@@ C.2 Observation Perturbation Details during the Testing Phase
We evaluate RDT and other baselines under two types of observation perturbations during the testing phase: Random and Action Diff perturbations, as described in prior works (Yang et al., 2022a; Zhang et al., 2020). Generally, offline RL algorithms are sensitive to observation perturbations, often resulting in a significant performance drop. The detailed implementations are as follows:
  • Random: We sample perturbed states within an $l_{\infty}$ ball of norm $\epsilon$ centering at the original state. Specifically, we create the perturbation set $\mathbb{B}_{d}(s,\epsilon)=\{\hat{s}:d(s,\hat{s})\leq\epsilon\}$ for state $s$, where $d(\cdot)$ is the $l_{\infty}$ norm, and sample one perturbed state to return to the agent.
  • Action Diff: This is an adversarial attack based on a pretrained IQL deterministic policy $\mu(s)$. We first sample 50 perturbed states within an $l_{\infty}$ ball of norm $\epsilon$ and then find the one that maximizes the difference in actions: $\max_{\hat{s}\in\mathbb{B}_{d}(s,\epsilon)}||\mu(s)-\mu(\hat{s})||^{2}$.
The parameter $\epsilon$ controls the scale of observation perturbations. We consider $\{0.0,0.1,0.3,0.5\}$ for $\epsilon$ in our experiments. In this setup, we first train offline RL algorithms under various data corruption settings, and then evaluate their performance in environments with observation perturbations.
@@ C.3 Implementation Details of RDT
We implement RDT and other baselines using the existing code base [^1] https://github.com/tinkoff-ai/CORL. Specifically. we build the network with 3 Transformer blocks, incorporating one MLP embedding layer for each key element: state, action, and return-to-go. We update the neural network using the AdamW optimizer, with a learning rate set at $1\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. The batch size is set to 64, with a sequence length of 20, to ensure effective and efficient training. To maintain stability during training, we adopt the state normalization as in Yang et al. (2024b). During the training phase, we train DeFog, DT and RDT for 100 epochs and other baselines with 1000 epochs. Each epoch is characterized by 1000 gradient steps. For evaluative purposes, we rollout each agent in environment across 10 trajectories, each with a maximum length of 1000, and we average the returns from these rollouts for comparison. We ensure the consistency and reliability of our reported results by averaging them over 4 unique random seeds. All experiments are conducted on P40 GPUs.
As for the hyperparameters of RDT, we configure the values for the embedding dropout probability as $p$, Gaussian Weighted coefficient $\beta_{a}$ and $\beta_{r}$, and iterative data correction threshold $\zeta$ through the sweeping mechanism. We start the iterative data correction at the 50th epoch for all tasks. We use iterative action correction exclusively for action corruption and iterative reward correction exclusively for reward corruption; therefore, only a single $\zeta$ value is shown in Table 3. It’s noteworthy that we do not conduct a hyperparameter search for individual datasets; instead, we select hyperparameters that apply to the entire task group. As such, RDT doesn’t require extensive hyperparameter tuning as previous methods. The exact hyperparameters used for the random corruption experiment are detailed in Table 3. We employ almost same hyperparameters for the adversarial corruption experiment, further attesting to the robustness and stability of RDT.
Table 3: Hyperparameters used for RDT under the random corruption.
@@
MuJoCo (10%) state 0.2 (1.0, 1.0) --
action 0.2 (1.0, 1.0) 6.0
reward 0.2 (1.0, 1.0) 6.0
Kitchen state 0.1 (30.0, 30.0) --
action 0.1 (30.0, 30.0) 5.0
reward 0.1 (30.0, 30.0) 6.0
@@
C.4 Detailed Information on Dataset
Since we conduct experiments on both the full and down-sampled datasets, we provide detailed information about the number of transitions and trajectories. As seen from Table 4, after down-sampling, the number of transitions remains of the same order of magnitude for most datasets, except for the Kitchen. Due to the small number of transitions in the original full “kitchen-complete” dataset, we opted not to perform down-sampling on the Kitchen.
Dataset halfcheetah hopper walker2d
# Transitions 202000 402000 302000
# Trajectories 202 2041 1093
Dataset halfcheetah(10%) hopper(10%) walker2d(10%)
# Transitions 20000 32921 27937
# Trajectories 20 204 109
Dataset kitchen-complete kitchen-partial kitchen-mixed
# Transitions 3680 136950 136950
# Trajectories 19 613 613
Dataset door(1%) hammer(1%) relocate(1%)
# Transitions 10000 10000 10000
# Trajectories 50 50 50
Table 4: Detailed information about the number of transitions and trajectories.